at what wavelength does a cavity at 5000 kelvin radiate most per unit wavelength??

Grade:12th pass

1 Answers

sonika p
6 years ago
according to stefan Boltzaman Law, 
(lambda)max * T = 2.9*10^-3
lambda = (2.9*10^-3)/5000 = 5.8* 10^-7 m